# Do Sell-side Analyst Reports Have Investment Value?

### v1

@@ 4 Conclusion
I also observe that return predictability is predominantly driven by reports conveying negative sentiment yet predicting positive future returns. This pattern suggests a systematic overreaction to negative news—such as weak quarterly earnings—even when a stock’s fundamentals remain strong. Further, the predictability is more pronounced for large, mature stocks and is particularly evident in reports authored by skilled and experienced analysts.
Next, I investigate which specific content within analyst reports contributes most to their investment value. Using a Shapley value decomposition, I show that Strategic Outlook content—especially statements revealing positive future fundamentals in the investment thesis and valuation models—dominates portfolio performance. Remarkably, a trading strategy based solely on strategic outlook content outperforms trading strategies based on full content.
This study provides the first direct evidence of the investment value embedded in analyst report text. These reports reveal analysts’ beliefs and expertise regarding particular stocks and convey valuable information to the market. More broadly, the findings illustrate the potential of artificial intelligence to enhance human judgment in financial contexts. Rather than replacing analyst insights, machine learning adds the most value when scaling and systematically extracting human-generated information from natural language.
@@ Appendix A Tables and Figures
This table reports the value-weighted decile portfolio performances sorted by the average of four sentiment measures of the past 12 months’ analyst reports. Contemporaneous Return denotes two-day DGTW characteristic-adjusted buy-and-hold abnormal returns following report release dates. Recommendation Revision denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same stock. Headline Sentiment represents the sentiment score of report headlines measured using a fine-tuned BERT model. Body Sentiment represents the average sentiment score of report body content measured using a fine-tuned BERT model. For each portfolio, I report several performance measures: excess return mean and standard deviation, Sharpe ratio, $\alpha$ relative to Fama and French (2015) five factors with momentum (six factors in total), and the corresponding t-statistics for $\alpha$. Portfolios are rebalanced monthly. The ‘H-L’ row represents a long-short strategy that takes a long position in the highest decile (High) and a short position in the lowest decile (Low). The sample period extends from January 2005 to June 2024.

### v2

@@ 4 Conclusion
I also observe that return predictability is predominantly driven by reports conveying negative sentiment yet predicting positive future returns. This pattern suggests a systematic overreaction to negative news, such as weak quarterly earnings, even when a stock’s fundamentals remain strong. Further, the predictability is more pronounced for large, mature stocks and is particularly evident in reports authored by skilled and experienced analysts.
Next, I investigate which specific content within analyst reports contributes most to their investment value. Using a Shapley value decomposition, I show that Strategic Outlook content, especially statements revealing positive future fundamentals in the investment thesis and valuation models, dominates portfolio performance. Remarkably, a trading strategy constructed univariately on strategic outlook content outperforms trading strategies based on full content.
This study provides the first direct evidence of the investment value embedded in analyst report text. These reports reveal analysts’ beliefs and expertise regarding particular stocks and convey valuable information to the market. More broadly, the findings illustrate the potential of artificial intelligence to enhance human judgment in financial contexts. Rather than replacing analyst insights, language models and machine learning add the most value when scaling and systematically extracting value-relevant information from natural language.
@@ Appendix A Tables and Figures
This table reports the value-weighted decile portfolio performances sorted by the average of four measures of the past 12 months’ analyst reports. Contemporaneous Return denotes two-day DGTW characteristic-adjusted buy-and-hold abnormal returns following report release dates. Recommendation Revision denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same stock. Headline Sentiment represents the sentiment score of report headlines measured using a fine-tuned BERT model. Body Sentiment represents the average sentiment score of report body content measured using a fine-tuned BERT model. For each portfolio, I report several performance measures: excess return mean and standard deviation, Sharpe ratio, $\alpha$ relative to Fama and French (2015) five factors with momentum (six factors in total), and the corresponding t-statistics for $\alpha$. Portfolios are rebalanced monthly. The ‘H-L’ row represents a long-short strategy that takes a long position in the highest decile (High) and a short position in the lowest decile (Low). The sample period extends from January 2005 to June 2024.